OpenStudy (unklerhaukus):

Solve the First Order Differential Equation:\[\frac{\text dy}{\text dx}-y\tan x=-y^2\sec x\]

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}=y\tan x-y^2\sec x\]

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}-y\tan x=-y^2\sec x\] \[\text {let }y=\frac 1v\] \[\frac{\text dy}{\text dx}=-\frac 1{v^2}\frac{\text dv}{\text dx}\] \[-\frac 1{v^2}\frac{\text dv}{\text dx}-\frac{\tan x}v=-\frac{\sec x}{v^2}\] \[\frac{\text dv}{\text dx}+v\tan x =\sec x\] \[v^\prime+v\tan x=\sec x\] \[R(x)=e^{\int \tan x\cdot\text dx}=e^{-\int\frac{-\sin x}{\cos x}\text dx}=e^{-\ln(\cos x)}=\sec x\] \[\left(v\sec x\right)^\prime=\sec^2 x\] \[v\sec x=\int \sec^2 x\text dx\] \[v\sec x=\tan x+c\] \[v=\sin x+c\cos x\] \[\frac 1y=\sin x+c\cos x\] \[y=\frac1{\sin x+c\cos x}\]